OpenStudy (anonymous):

The legs of an isosceles right triangle are 11 cm long. Find the length of the hypotenuse. Round to the nearest tenth if necessary

OpenStudy (anonymous):

11^2+11^2=x^2 x^2=121+121=242 x=11*sqrt2

OpenStudy (anonymous):

Pythagoras' theorem tells us that \[a^2+b^2=c^2\]We know a and b are 11, and we want to find c. Substitute a and b into the equation and rearrange to solve for c.

OpenStudy (mertsj):

Use what you know about the special right triangles. The leg multiplied by sqrt2 will give you the hypotenuse of an isosceles right triangle. So multiply 11 by 1.414